Sorry if this is the wrong place for this question.
If there are 2 apples and every year the percentage of apples increase by 2%. How many apples are there in 6000 years?

Answer by lazi(15)

starting out with 2 apples
every year the amount of apples increase by 2%
how many apples are there in 6000 years
2+1[2(.02)]=2.04
2+2[2(.02)]=2.08
and so on to 6000 years
2+6000[2(.02)]=242
the beginning 2 is your starting out amount.
6000 is your years, and 2(.02) is the amount increase per year.
another way of saying this is that in 6000 years the % increase is (2% * 6000) = 12000% = 12000/100
2+2*12000/100= 2+240=242